# Complete Characterizations of Well-Posedness in Parametric Composite Optimization

### v2

@@ 5 Characterizations of Lipschitzian Stability for KKT Systems
Assuming the contrary, for any $k$ sufficiently take a sequence $\delta^{k}\in\frac{1}{k}\mathbb{B}_{\mathcal{R}^{n}}$ and find $(x^{k},u^{k})\in S_{\operatorname{KKT}}(\delta^{k},0)$ with $x^{k}\rightarrow\bar{x}$ and $w^{k}\neq 0$ with $\|w^{k}\|=1$ such that
@@
(ii) SOQC (4.1) holds together with the second-order kernel condition
@@
If $\bar{x}$ is a local minimizer of problem (1.1), $(A1)$ is fulfilled for $g$ at $F(\bar{x})$, the above are equivalent to each of the properties:
@@
(vi) SOQC (4.1) holds and $\bar{x}$ is a fully stable local minimizer of problem (1.2).
@@
Furthermore, in the case where $g$ is $\mathcal{C}^{2}$-cone reducible at $F(\bar{x})$, SOQC (4.1) in
(iii), (iv), and (vi) can be equivalently replaced by the constraint nondegeneracy condition (3.4).
@@
[…]
@@
Hence the inclusion $(0,0)\in D\Phi((x,u),(\delta,0))(\Delta\xi^{\prime},\Delta\eta^{\prime})$ yields $\nabla_{xx}^{2}L(x,u)\Delta\xi^{\prime}+\nabla F(x)^{T}\Delta\eta^{\prime}=0$. It follows from [MordukhovichTangWang-tilt-2025, Proposition 3.2] that $g$ is twice epi-differentiable at the points in question. This leads us by using (2.11), (2.12), and [Rockafellar1998, Theorem 13.57] that
@@
and thus $(0,0)\in D^{*}\Phi((x,u),(\delta,0))(\Delta\xi^{\prime},-\Delta\eta^{\prime})$. Employing the characterization of Lemma 2.1, we deduce from the Aubin property of $S_{\operatorname{KKT}}$ at $(\delta,0)$ for $(x,u)$ that $\Delta\xi^{\prime}=0$ and $\Delta\eta^{\prime}=0$. Then the Levy-Rockafellar criterion in Lemma 2.2 tells us that $S_{\operatorname{KKT}}$ enjoys the isolated calmness property at $(\delta,0)$ for $(x,u)$ and $S_{\operatorname{KKT}}(\delta,0)$ is a singleton. Let $\delta=0$, Lemma 5.1, valid under (A1), confirms that SOQC (4.1) and SOSC (5.1) hold for $(\bar{x},\bar{u})$. Consequently, this guarantees that the second-order growth condition is satisfied for problem (1.1) at $\bar{x}$ (see, e.g., [Mohammadi2020, Theorem 6.1] ) and RCQ (3.1) holds at $\bar{x}$. It is straightforward to verify that the second-order growth condition for problem (4.7) also holds at the point $(\bar{x},g(F(\bar{x})))$. Applying [Bonnans2000, Proposition 4.32], we conclude that the solution set of problem (5.2) with the parameter $a=\delta$ is nonempty. This result implies that the solution set of problem (1.2) with parameters $(a,b)=(\delta,0)$ is also nonempty. Given any $\delta$ with its norm sufficiently small, let $x^{\delta}$ be a local minimizer of problem (1.2) with $(a,b)=(\delta,0)$. By [Bonnans2000, Proposition 4.32], $x^{\delta}$ is sufficiently close to $\bar{x}$ and RCQ holds at $x^{\delta}$. There exists a unique $u^{\delta}\in\Lambda(x^{\delta},\delta,0)$ with $\Lambda(x^{\delta},\delta,0)=\{u^{\delta}\}$ and $S_{\operatorname{KKT}}(\delta,0)=\Phi^{-1}(\delta,0)=\{(x^{\delta},u^{\delta})\}$ both singletons. We conclude that for any $(x,u)\in S_{\operatorname{KKT}}(\delta,0)$, $x$ is a local minimizer of problem (1.2) with $(a,b)=(\delta,0)$ and $S_{\operatorname{KKT}}$ has the isolated calmness property at $(\delta,0)$ for $(x,u)$. Applying Lemma 5.1, it follows that SOQC (4.1) and SOSC (5.1) hold for any $(x,u)\in\Phi^{-1}(\delta,0)=\{(x^{\delta},u^{\delta})\}$ with $\delta\in\mathcal{V}$ and $(x,u)\in\mathcal{U}$. Deduce from [MordukhovichTangWang-tilt-2025, Theorem 4.1] the existence of $\kappa>0$ such that $\bar{x}$ is a tilt-stable local minimizer of $\varphi$ with modulus $\kappa$, which justifies assertion (iv).
@@
$\min_{z\in D^{*}(\partial g)(F(\bar{x}),\bar{u})(\nabla F(\bar{x})d)}\left\langle\nabla F(\bar{x})d,z\right\rangle$ $\leq\left\langle\nabla F(\bar{x})d,v\right\rangle=-\left\langle\nabla_{xx}^{2}L(\bar{x},\bar{u})d,d\right\rangle,$
which tells us in turn that the chosen vector $d\neq 0$ satisfies the inequality
$\left\langle\nabla^{2}_{xx}L(\bar{x},\bar{u})d,d\right\rangle+\min_{z\in D^{*}(\partial g)(F(\bar{x}),\bar{u})(\nabla F(\bar{x})d)}\left\langle\nabla F(\bar{x})d,z\right\rangle\leq 0.$
@@
As mentioned above, assumption (A1) is valid for all $\mathcal{C}^{2}$-cone reducible functions. Consequently, all the conditions in Theorem 5.1 are readily equivalent when $g$ is the indicator function of the second-order cone, of the positive semidefinite cone, the nuclear norm function, the matrix spectral norm function, and in other settings important in applications.
The next lemma, important for its own sake, provides a useful characterization of strong regularity for the KKT system (1.3) via the strict graphical derivative (2.8).
Lemma 5.4.
Let $(\bar{x},\bar{u})$ be a solution to the KKT system (1.3) with $(a,b)=(0,0)$. Then the pair $(\bar{x},\bar{u})$ is a strongly regular point of (1.3) if and only if we have
$\operatorname{ker}(\nabla F(\bar{x})^{T})\cap D_{*}(\partial g)(F(\bar{x}),\bar{u})(0)=\{0\},$ (5.9)
$\left\{d\,|\,0\in\nabla^{2}_{xx}L(\bar{x},\bar{u})d+\nabla F(\bar{x})^{T}D_{*}(\partial g)(F(\bar{x}),\bar{u})(\nabla F(\bar{x})d)\right\}=\{0\}.$ (5.10)
Proof.
We know from [BenkoRockafellar2024, Theorem 4.2] that the strong regularity of $(\bar{x},\bar{u})$ is equivalent to
$\left[\begin{array}[]{c}0\\ 0\end{array}\right]\in\left[\begin{array}[]{c}\nabla_{xx}^{2}L(\bar{x},\bar{u})\Delta\xi+\nabla F(\bar{x})^{T}\Delta\eta\\ -\Delta\eta+D_{*}(\partial g)(F(\bar{x}),\bar{u})(\nabla F(\bar{x})\Delta\xi)\end{array}\right]\,\Longrightarrow\,\left[\begin{array}[]{c}\Delta\xi\\ \Delta\eta\end{array}\right]=\left[\begin{array}[]{c}0\\ 0\end{array}\right],$ (5.19)
which verifies (5.10) and tells us that if $(0,\Delta\eta)$ is a solution to the left-hand side system in (5.19), then $\Delta\eta=0$, i.e., (5.9) holds. Substituting $\Delta\eta$ from the second equation into the first one yields (5.10). Conversely, if $(\Delta\xi,\Delta\eta)$ is a solution to the left-hand side system in (5.19), then $\Delta\xi=0$ by (5.10). Thus it follows from (5.9) that $\Delta\eta=0$, which completes the proof. ∎
As a consequence of Theorem 5.1 with the help of Lemma 5.4, we get the equivalence between the Aubin property and the strong regularity for a specific class of generalized equations.
Corollary 5.1.
Let $\bar{x}$ be a local minimizer of problem (1.1), where $g$ is $\mathcal{C}^{2}$-cone reducible and the Jacobian $\nabla F(\bar{x})$ is of full rank. Then conditions (5.6) and (5.10) are equivalent. In the case where $F(x):=x$, we have that the Aubin property of the solution mapping to the generalized equation
$0\in\Psi(x):=\nabla h(x)+\partial g(x)$ (5.20)
at the origin for $\bar{x}$ is equivalent to the strong regularity of $\bar{x}$ for $\Psi^{-1}$ of (5.20).
Proof.
When $\nabla F(\bar{x})$ is of full rank, SOQC (4.1) and condition (5.9) are automatically satisfied. Under the $\mathcal{C}^{2}$-cone reducibility of $g$, Theorem 5.1 and Lemma 5.4 imply that (5.6) is equivalent to (5.10). Then the equivalence claimed in the corollary follows from Theorem 5.1 due to Lemmas 2.1 and 2.3 combined with (2.14) and (2.15). ∎
Remark 5.1.
The equivalence result related to Corollary 5.1 can be deduced from [sun25, Theorem 4] devoted to the generalized equation defined by $\Psi(x):=f(x)+N_{\Omega}(x)$, where $\Omega\subset\mathcal{R}^{n}$ is a ${\cal C}^{2}$-cone reducible set. Using the degree theory, [sun25, Theorem 4] establishes the equivalence between the Aubin property and strong regularity of $\Psi^{-1}$ at the reference point. The latter result can be specified for optimization problems of the form $\min_{x\in\mathcal{R}^{n}}h(x)+\delta_{\Omega}(x)$, where the nondegeneracy condition is automatically satisfied. We see that Corollary 5.1 and [sun25, Theorem 4], derived by completely different methods, overlap. In particular, when $g(x)=\delta_{\Omega}(x)$ with $\Omega$ being a convex $\mathcal{C}^{2}$-cone reducible set, Corollary 5.1 gives us the same conclusion as [sun25, Theorem 4] for the optimization problems under consideration.
@@
Let $\bar{x}$ be a local minimizer of problem (1.1) with $\bar{u}\in\Lambda(\bar{x},0,0)$ as a corresponding Lagrange multiplier, where $g$ satisfies (A1) and (A2) at $F(\bar{x})$ for $\bar{u}$. If SOQC (4.1) and SSOSC (3.10) hold at $\bar{x}$, then all the matrices from the generalized Jacobian $\mathcal{J}Z(\bar{x},\bar{u})$ are nonsingular.
@@
Combining this with SSOSC (3.10) tells us that $\Delta x=0$. Substituting $\Delta x=0$ into (5.24) and (5.25), we get $\Delta u\in\operatorname{ker}(\nabla F(\bar{x})^{T})\cap\operatorname{ker}U$. Applying [Tangwang2024, Corollary 4.1], (2.12) and (2.14) leads us to
@@
and thus ensures that SOQC (4.1) yields $\Delta u=0$, which completes the proof. ∎
@@
Let $\bar{x}$ be a local minimizer of problem (1.1) with $\bar{u}\in\Lambda(\bar{x},0,0)$ as a corresponding Lagrange multiplier, and let $g$ satisfy both assumptions (A1), (A2) at $F(\bar{x})$ for $\bar{u}$. Then we have the equivalent assertions:
(i) The mapping $S_{\operatorname{KKT}}$ has the Aubin property at the origin for $(\bar{x},\bar{u})$.
@@
(vi) SOQC (4.1) holds and $\bar{x}$ is a fully stable local minimizer of problem (1.2).
@@ 6 Concluding Remarks and Future Research
This paper establishes equivalent relationships between various well-posedness properties of general composite optimization problems and associated KKT systems with providing comprehensive second-order characterizations of these properties via pointbased constructions of second-order generalized differentiation. In particular, we resolve at the alternative a long-standing question about the equivalence between the Aubin property and strong regularity of solution mappings to KKT systems, as well as the tilt stability of local minimizers, in the ${\cal C}^{2}$-cone reducible setting of composite optimization with parabolically regular data. Such an equivalence has been established earlier for important special classes of problems in constrained optimization. We also show, among other results, that these properties and their equivalents are completely characterized by the simultaneous fulfillment of certain constraint nondegeneracy and robust second-order subdifferential conditions expressed precisely at the reference solution and satisfying comprehensive calculus rules. The obtained results are of a fundamental nature, while being instrumental for stability and sensitivity analysis of optimization-related problems, as well as for the design and justification of numerical algorithms in constrained optimization.
Our future research will be conducted in various directions. Among them, we plan to reveal—besides the broad settings discussed above—other classes of composite optimization problems satisfying assumption (A2) with explicit computations of the second-order variational function $\Gamma_{f}$. Recalling that the strong regularity, Lipschitzian stability of KKT systems, and tilt stability of local minimizers play a crucial role in justifying numerical algorithms of constrained optimization, our attention will be paid to the implementation of these ideas and the obtained results to new classes of composite optimization problems. Moreover, we intend to investigate some of the well-posedness properties studied in this paper without imposing nondegeneracy conditions.
Acknowledgments. The authors are grateful to Helmut Gfrerer for valuable remarks, especially concerning Lemma 5.1 and its applications.

### v3

@@ 5 Characterizations of Lipschitzian Stability for KKT Systems
Assuming the contrary, for any $k$ sufficiently large take a sequence $\delta^{k}\in\frac{1}{k}\mathbb{B}_{\mathcal{R}^{n}}$ and find $(x^{k},u^{k})\in S_{\operatorname{KKT}}(\delta^{k},0)$ with $x^{k}\rightarrow\bar{x}$ and $w^{k}\neq 0$ with $\|w^{k}\|=1$ such that
@@
(ii) SOQC (4.1) holds together with the second-order kernel condition at $\bar{x}$ for $\bar{u}$
@@
If $\bar{x}$ is a local minimizer of (1.1) and $(A1)$ is fulfilled for $g$ at $F(\bar{x})$, the following are equivalent:
(i’) The KKT solution mapping $S_{\operatorname{KKT}}$ has the Aubin property at the origin for $(\bar{x},\bar{u})$. There exist neighborhoods $\mathcal{U}$ of $(\bar{x},\bar{u})$ and $\mathcal{V}$ of the origin such that for $\delta\in\mathcal{V}$, the point $x\in\{x\;|\;(x,u)\in S_{\operatorname{KKT}}(\delta,0)\cap\mathcal{U}\}$ is a local minimizer of problem (1.2) with $(a,b)=(\delta,0)$.
@@
(vi) The constraint nondegeneracy condition (3.4) holds at $\bar{x}$ for $\bar{u}$ and $\bar{x}$ is a fully stable local minimizer of problem (1.2).
@@
that $\Delta\eta=0$. The reverse implication is easy. Indeed, the second-order kernel condition gives us $\Delta\xi=0$ in (5.8), which ensures by SOQC at $x$ for $u$ that $\Delta\eta=0$. Then the Aubin property follows directly from the characterization in Lemma 2.1. Letting $(x,u):=(\bar{x},\bar{u})$ and $\delta=0$ yields the equivalence between the Aubin property of $S_{\operatorname{KKT}}$ at the origin for $(\bar{x},\bar{u})$ and SOQC (4.1) combined with the second-order kernel condition (5.6).
@@
Hence the inclusion $(0,0)\in D\Phi((x,u),(\delta,0))(\Delta\xi^{\prime},\Delta\eta^{\prime})$ yields $\nabla_{xx}^{2}L(x,u)\Delta\xi^{\prime}+\nabla F(x)^{T}\Delta\eta^{\prime}=0$. It follows from [29, Proposition 3.2] that $g$ is twice epi-differentiable at the points in question. This leads us by using (2.11), (2.12), and [39, Theorem 13.57] that
@@
and thus $(0,0)\in D^{*}\Phi((x,u),(\delta,0))(\Delta\xi^{\prime},-\Delta\eta^{\prime})$. Employing the characterization of Lemma 2.1, we deduce from the Aubin property of $S_{\operatorname{KKT}}$ at $(\delta,0)$ for $(x,u)$ that $\Delta\xi^{\prime}=0$ and $\Delta\eta^{\prime}=0$. Then the Levy-Rockafellar criterion in Lemma 2.2 tells us that $S_{\operatorname{KKT}}$ enjoys the isolated calmness property at $(\delta,0)$ for $(x,u)$. Let $\delta=0$, Lemma 5.1, valid under (A1), confirms that SRCQ (3.3) and SOSC (5.1) hold for $(\bar{x},\bar{u})$. Consequently, this guarantees that the second-order growth condition is satisfied for problem (1.1) at $\bar{x}$ (see, e.g., [23, Theorem 6.1] ) and RCQ (3.1) holds at $\bar{x}$. It is straightforward to verify that the second-order growth condition for problem (4.8) also holds at the point $(\bar{x},g(F(\bar{x})))$. Applying [4, Proposition 4.32], we conclude that the solution set of problem (5.2) with the parameter $a=\delta$ is nonempty. This result implies that the solution set of problem (1.2) with parameters $(a,b)=(\delta,0)$ is also nonempty. Given any $\delta$ with its norm sufficiently small, let $x^{\delta}$ be a local minimizer of problem (1.2) with $(a,b)=(\delta,0)$. By [4, Proposition 4.32], $x^{\delta}$ is sufficiently close to $\bar{x}$. The robustness property of RCQ implies RCQ holds at $x^{\delta}$ and there exists a unique $u^{\delta}\in\Lambda(x^{\delta},\delta,0)$ with $\Lambda(x^{\delta},\delta,0)=\{u^{\delta}\}$. We conclude that for any $(x,u)\in S_{\operatorname{KKT}}(\delta,0)$, $x$ is a local minimizer of problem (1.2) with $(a,b)=(\delta,0)$ and $S_{\operatorname{KKT}}$ has the isolated calmness property at $(\delta,0)$ for $(x,u)$. Applying Lemma 5.1, it follows that SOSC (5.1) holds for any $(x,u)\in\Phi^{-1}(\delta,0)$ with $\delta\in\mathcal{V}$ and $(x,u)\in\mathcal{U}$. Deduce from [29, Theorem 4.1] the existence of $\kappa>0$ such that $\bar{x}$ is a tilt-stable local minimizer of $\varphi$ with modulus $\kappa$, which justifies assertion (iv).
@@
$\langle d^{k},w^{k}\rangle$ $=\langle\nabla_{xx}^{2}L(x^{k},u^{k})d^{k},d^{k}\rangle+d^{2}g(F(x^{k}),u^{k})(\nabla F(x^{k})d^{k})$
$=\langle\nabla_{xx}^{2}L(x^{k},u^{k})d^{k},d^{k}\rangle+\langle\nabla F(x^{k})d^{k},v^{k}\rangle.$
By [39, Theorem 13.57] and [29, Proposition 3.2], taking a subsequence if necessary and letting $k\rightarrow\infty$, the coderivative robustness tells us that there exist $v\in D^{*}(\partial g)(F(\bar{x}),\bar{u})(\nabla F(\bar{x})d)$ with
$\langle d^{k},w^{k}\rangle\rightarrow\langle\nabla_{xx}^{2}L(\bar{x},\bar{u})d,d\rangle+\langle\nabla F(\bar{x})d,v\rangle\leq 0,$
which contradicts the second-order subdifferential condition (5.3) and thus verifies (iv).
@@
Further, it follows from Lemma 4.2, (iv), Proposition 2.1, and [29, Theorem 4.1] that (i) is equivalent to the Aubin property of the KKT mapping for the auxiliary problem (4.8) at the origin for $((\bar{x},g(F(\bar{x}))),(\bar{u},-1))$. Therefore, (i’) is also equivalent to SOQC (4.1) combined with the second-order subdifferential condition for problem (4.8). The strong regularity of $((\bar{x},g(F(\bar{x}))),(\bar{u},-1))$ for the KKT system associated with problem (4.8) directly implies the fulfillment of this property of $(\bar{x},\bar{u})$ for the KKT system (1.3), which consequently guarantees the Aubin property of the solution mapping. Moreover, if $(\bar{x},g(F(\bar{x})))$ is a fully stable minimizer of problem (4.8), then $\bar{x}$ is fully stable for problem (1.1), since the full stability yields the tilt-stability of $\bar{x}$. Remembering that $\bar{x}$ is a local minimizer of $\varphi$ allows us to deduce from [42, Lemma 4.2], (2.11), and (2.13) that
$\begin{array}[]{ll}&\mathcal{C}(F(\bar{x}),g(F(\bar{x})))=\{(d,\beta)\;|\;dg(F(\bar{x}))(d)=\langle\bar{u},d\rangle,\;\beta=0\},\\ &\operatorname{dom}D^{*}\mathcal{N}_{\operatorname{epi}g}((F(\bar{x}),g(F(\bar{x}))),(\bar{u},-1))\subset\{(d,0)\;|\;d\in\operatorname{aff}\{y\;|\;dg(F(\bar{x}))(y)=\langle\bar{u},y\rangle\}\}.\end{array}$
@@
Let $(\bar{x},\bar{u})$ be a solution to the KKT system (1.3) with $(a,b)=(0,0)$, where $g$ satisfies (A1) and (A2) at $F(\bar{x})$ for $\bar{u}$. If SOQC (4.1) and SSOSC (3.10) hold at $\bar{x}$, then all the matrices from the generalized Jacobian $\mathcal{J}Z(\bar{x},\bar{u})$ are nonsingular.
@@
Combining this with SSOSC (3.10) yields $\Delta x=0$. Substituting $\Delta x=0$ into (5.12) and (5.13), we get $\Delta u\in\operatorname{ker}(\nabla F(\bar{x})^{T})\cap\operatorname{ker}U$. Applying [42, Corollary 4.1], (2.12), and (2.14) leads us to
@@
and thus ensures that SOQC (4.1) implies $\Delta u=0$, which completes the proof. ∎
@@
Let $\bar{x}$ be a local minimizer of problem (1.1) with $\bar{u}\in\Lambda(\bar{x},0,0)$ as a corresponding Lagrange multiplier, and let $g$ satisfy both assumptions (A1), (A2) at $F(\bar{x})$ for $\bar{u}$. Then we have the following equivalent assertions:
(i) The mapping $S_{\operatorname{KKT}}$ has the Aubin property at the origin for $(\bar{x},\bar{u})$. There exist neighborhoods $\mathcal{U}$ of $(\bar{x},\bar{u})$ and $\mathcal{V}$ of the origin such that for $\delta\in\mathcal{V}$, the point $x\in\{x\;|\;(x,u)\in S_{\operatorname{KKT}}(\delta,0)\cap\mathcal{U}\}$ is a local minimizer of problem (1.2) with $(a,b)=(\delta,0)$.
@@
(i) The KKT solution mapping $S_{\operatorname{KKT}}$ has the Aubin property at the origin for $(\bar{x},\bar{u})$. There exist neighborhoods $\mathcal{U}$ of $(\bar{x},\bar{u})$ and $\mathcal{V}$ of the origin such that for $\delta\in\mathcal{V}$, the point $x\in\{x\;|\;(x,u)\in S_{\operatorname{KKT}}(\delta,0)\cap\mathcal{U}\}$ is a local minimizer of problem (1.2) with $(a,b)=(\delta,0)$.
@@
(vi) The constraint nondegeneracy condition (3.4) holds at $\bar{x}$ for $\bar{u}$ and $\bar{x}$ is a fully stable local minimizer of problem (1.2).
@@ 6 Concluding Remarks and Future Research
This paper establishes equivalent relationships between various well-posedness properties of general composite optimization problems and associated KKT systems with providing second-order characterizations of these properties via pointbased constructions of second-order generalized differentiation. In particular, we establish close relationships between the Aubin property and strong regularity of solution mappings to KKT systems, as well as the tilt stability of local minimizers, in the ${\cal C}^{2}$-cone reducible setting of composite optimization with parabolically regular data. It is also shown, among other results, that these properties and their equivalents are characterized by the simultaneous fulfillment of certain constraint nondegeneracy and robust second-order subdifferential conditions expressed precisely at the reference solution and satisfying comprehensive calculus rules. The obtained results are of a fundamental nature, while being instrumental for stability and sensitivity analysis of optimization-related problems, as well as for the design and justification of numerical algorithms in constrained optimization.
Our future research will be conducted in various directions. Among them, we plan to reveal—besides the broad settings discussed above—other classes of composite optimization problems satisfying assumption (A2) with explicit computations of the second-order variational function $\Gamma_{f}$. We also intend to fully investigate the possibility to eliminate the assumption on the KKT solution mapping $S_{KKT}$ imposed in Theorem 5.1(i’), Theorem 5.2(i), and Corollary 5.1(i). Recalling that the strong regularity, Lipschitzian stability of KKT systems, and tilt stability of local minimizers play a crucial role in justifying numerical algorithms of constrained optimization, our attention will be paid to the implementation of these ideas and the obtained results to new classes of composite optimization problems. Moreover, we intend to investigate some of the well-posedness properties studied in this paper without imposing nondegeneracy conditions.
Acknowledgments. The authors are gratefully indebted to Helmut Gfrerer for observing serious inaccuracies in Section 5 and very useful suggestions that allowed us to significantly improve the original presentation.
Disclosure of interest
Research of Boris Mordukhovich was partly supported by the US National Science Foundation under grant DMS-2204519 and by the Australian Research Council under Discovery Project DP250101112. Research of Peipei Tang was supported by Zhejiang Provincial Natural Science Foundation of China under Grant No. LMS26A010022. Research of Chengjing Wang was partly supported by the National Natural Science Foundation of China under Grant U21A20169, by the Zhejiang Provincial Natural Science Foundation of China under Grant LTGY23H240002. The authors report there are no competing interests to declare.
Data Availability
Data sharing is not applicable to this paper as no new data were created or analyzed.